3:48 Cm F CC=70) + (T(n) = T(n − 1) +n T(1) = 10
Answer #1

3. TCM = T(n-1) + and T(I) = 10. 1 T(n = T(3-1) en = T (0-2) + (-1) an. = T(n-3) + (0-2 (0-1) an. = T(N-K) & (n-K+1) of (n-k+